Use the Following Information for Questions 1 and 2.

A firm currently has earnings of $2 per share and pays out 30% of earnings as dividends on its common stock. The after tax return on equity is 15%. The investor requires a 17% return.

1.What is the estimated growth rate of earnings and dividends?

2.Using the constant growth model, what is the intrinsic value of the common stock?$

3. Three years from now you predict that a common stock will sell for $30. The stock pays dividends of $1.50, $1.75 and $2.00 for the next three years. If you require a 12 percent rate of return, what is the current value of the stock?

4.The dividend on a common stock just paid this year was $2.25 while the dividend paid 10 years ago was $1.05. The current dividend payout ratio is 40% and the investor's required return is 10%. What is the current P/E ratio for this stock?

5.You estimate that a stock's dividend will grow at a rate of 12% for two years and then slow to 3% forever and ever and ever and ever... You require a 12% rate of return and the current dividend is $2 per share. What is the intrinsic value of the stock?

6.Tiger Inc. retains 70% of its earnings and pays out the rest in dividends. Net income this year was $18 million, while stockholder's equity was $120 million. Current dividends are $4 per share. The T-bill rate is 4% and the return on the market index is expected to be 11%. Tiger Inc. has a β of 3.25. What is the price of this stock?
Answer #1

1) growth, g = ROE x Retention Ratio = 15% x (1 - 30%) = 10.5%

2) Value = D0 x (1 + g) / (r - g) = 2 x 30% x (1 + 10.5%) / (17% - 10.5%) = $10.20

3) Value = D1 / (1 + r) + D2 / (1 + r)^2 + (D3 + P3) / (1 + r)^3

= 1.50 / 1.12 + 1.75 / 1.12^2 + (2 + 30) / 1.12^3

= $25.51

4) Growth, g = (2.25 / 1.05)^(1/10) - 1 = 7.42%

P/E = Payout ratio x (1 + g) / (r - g) = 40% x (1 + 7.42%) / (10% - 7.42%) = 16.64